# FAILURE OF LICHNEROWICZ-TYPE RESULT IN PARABOLIC GEOMETRIES OF REAL RANK AT LEAST 33

### v3

@@ 2.7 Curvature Trees and Harmonic Seeds
Given $\Omega\in\ker\Delta$, let $K_{\Omega}=\mathrm{Stab}_{G_{0}}(\Omega)$ and let $\mathfrak{k}_{\Omega}\leq\mathfrak{g}_{0}$ be its Lie algebra. A form $\Omega\in(\ker\Delta)_{+}$ is said to have the Kruglikov-The property if
@@
If $\Omega$ has the Kruglikov-The property, it is said to be a harmonic seed if there exists a model geometry $(J_{\Omega},K_{\Omega})$ and an isomorphism of $K_{\Omega}$-representations $\psi:\mathfrak{j}_{\Omega}\to\mathfrak{g}_{-}\oplus\mathfrak{k}_{\Omega}\leq\mathfrak{g}$ such that $\mathfrak{j}_{\Omega}$ is the Lie algebra of $J_{\Omega},$ $\psi|_{\mathfrak{k}_{\Omega}}=1_{\mathfrak{k}_{\Omega}}$ and $J_{\Omega}/K_{\Omega}$ is simply connected. Let $\omega_{J_{\Omega}}$ and $\omega_{P}$ be the Maurer-Cartan forms of the Lie groups $J_{\Omega}$ and $P$. Then the Cartan geometry $\mathcal{G}_{\Omega}:=J_{\Omega}\times_{K_{\Omega}}P$ modeled on $(G,P)$ over $J_{\Omega}/K_{\Omega}$ with Cartan form defined by
@@
is called the curvature tree grown from $\Omega.$
For $j\in J_{\Omega},$ let $L_{j}:\mathcal{G}\to\mathcal{G}$ denote left-action by $j.$ This transformation is right $P$-equivariant. Also,
@@
so $J_{\Omega}\leq\mathrm{Aut}(\mathcal{G},\omega_{\Omega}).$ The following result is Theorem 3.4 in [6].
@@
Denote by $\mathfrak{b}_{-}$ the nilpotent subalgebra of $\mathfrak{g}$ generated by the negative restricted root spaces. If $\Omega\in(\ker\Delta)_{+}$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}$ then $\Omega$ is a harmonic seed.
@@ 3 Lowest Weights
This section develops results enabling us to compute lowest weight vectors in the module of harmonic curvature forms. With these results in hand, we can prove Theorem 3.10 and Theorem 3.12, which are useful for constructing harmonic seeds and compact quotients of curvature trees, respectively. Let $\mathfrak{g}$ be real semisimple. We keep in mind the identification $H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})\cong\ker\Delta\leq C^{2}(\mathfrak{g}_{-},\mathfrak{g})$. Given $\beta,\gamma\in\hat{\Delta}^{+}(\mathfrak{p}_{+})$, $\zeta\in\hat{\Delta}$, define a subspace $V_{\beta,\gamma,\zeta}:=(\mathfrak{g}_{\beta})_{\flat}\wedge(\mathfrak{g}_{\gamma})_{\flat}\otimes\mathfrak{g}_{\zeta}\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$.
@@
Let $V\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$ be a $\mathfrak{g}_{0}$-irreducible representation. Suppose there is a $\mathfrak{g}_{0}$ lowest weight vector $v\in V$ contained in $V_{\beta,\gamma,\zeta}.$ Then every $\mathfrak{g}_{0}$ lowest weight vector in $V$ is contained in $V_{\beta,\gamma,\zeta}$.
@@
We claim that every lowest weight vector in $V$ is in the $\mathfrak{z}(\mathfrak{g}_{0})$ module generated by $v,$ from which the result follows. Let $w\in V$ be another lowest weight vector. Let $\mathfrak{g}_{0}^{\geq 0}:=\mathfrak{z}(\mathfrak{g}_{0})\oplus\bigoplus_{\alpha\in\Delta^{+}(\mathfrak{g}_{0})}\mathfrak{g}_{\alpha}.$ Since $v$ is a lowest weight element, an inductive argument shows that $V$, the $\mathfrak{g}_{0}$ module generated by $v$, is in fact the $\mathfrak{g}_{0}^{\geq 0}$ module generated by $v$. The lowest restricted weight of an irreducible representation is unique, so $v$ and $w$ both have the same restricted weight. Because $w$ is in the $\mathfrak{g}_{0}^{\geq 0}$ module generated by $v$ but has the same restricted weight, it must be in the $\mathfrak{z}(\mathfrak{g}_{0})$ module generated by $v.$
@@
Vectors in $V_{\beta,\gamma,\zeta}$ have weight $\beta+\gamma+\zeta.$
@@
We have $(\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g})^{\mathbb{C}}\cong\bigwedge^{2}(\mathfrak{g}^{\mathbb{C}}_{-})^{*}\otimes\mathfrak{g}^{\mathbb{C}}$ and this isomorphism restricts to the isomorphism of Proposition 2.8(a) on harmonic elements.
Proposition 3.4.
Suppose $\beta,\gamma,\zeta$ are the restrictions of $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ to $\mathfrak{a}.$ Then $V_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}\leq(V_{\beta,\gamma,\zeta})^{\mathbb{C}}$.
@@
[…]
$\mathfrak{g}^{\mathbb{C}}=\mathfrak{z}(\mathfrak{a})^{\mathbb{C}}\oplus\bigoplus_{\alpha\in\hat{\Delta}}(\mathfrak{g}_{\alpha})^{\mathbb{C}},$
$(\mathfrak{g}_{\alpha})^{\mathbb{C}}$ must be the entire $\alpha$-eigenspace for the action of $\mathfrak{a}.$ We also have
$H,X]=\widetilde{\alpha}(H)X=\alpha(H)X$
for $H\in\mathfrak{a}$ and $X\in(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\alpha}}$. Therefore $(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\alpha}}\leq(\mathfrak{g}_{\alpha})^{\mathbb{C}}.$ The conclusion follows. ∎
@@
Suppose $V\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$ is a $\mathfrak{g}_{0}$ subrepresentation, and $\widetilde{\Omega}\in V^{\mathbb{C}}$ is a $\mathfrak{g}_{0}^{\mathbb{C}}$ weight vector such that $\widetilde{\Omega}\in V_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}$. Then the real and imaginary parts of $\widetilde{\Omega}$ in $V$, if nonzero, are $\mathfrak{g}_{0}$ weight vectors in $V_{\beta,\gamma,\zeta}$, where $\beta,\gamma,\zeta$ are the restrictions of $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ to $\mathfrak{a}.$ In addition, if $\widetilde{\Omega}$ is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector, then its real and imaginary parts, if nonzero, are $\mathfrak{g}_{0}$ lowest weight vectors.
@@
Without loss of generality, we prove the result for the real part. Define $\Omega,\Omega^{\prime}\in V$ as the unique vectors such that $\tilde{\Omega}=\Omega+i\Omega^{\prime}.$ Since $\tilde{\Omega}\in V_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}\leq(V_{\beta,\gamma,\zeta})^{\mathbb{C}},$ we have $\Omega,\Omega^{\prime}\in V_{\beta,\gamma,\zeta}.$ By assumption, $\Omega\neq 0$. Because $\tilde{\Omega}$ is a weight vector for $\mathfrak{g}_{0}^{\mathbb{C}},$ it is scaled by real values under the adjoint action of $\mathfrak{a},$ so $\Omega$ is a weight vector for $\mathfrak{g}_{0}.$
Now suppose $\widetilde{\Omega}$ is a lowest weight. Let $\alpha$ be some negative $\mathfrak{g}_{0}$ restricted root. Then $((\mathfrak{g}_{0})_{\alpha})^{\mathbb{C}}$ is a direct sum of negative rootspaces of $\mathfrak{g}^{\mathbb{C}}.$ These rootspaces annihilate $\tilde{\Omega}$ by assumption, so $(\mathfrak{g}_{0})_{\alpha}$ annihilates $\tilde{\Omega}$, so $(\mathfrak{g}_{0})_{\alpha}$ annihilates $\Omega.$ It follows that $\Omega$ is a lowest weight vector of $\mathfrak{g}_{0}$. ∎
As discussed in Section 2.4, harmonic curvature of regular, normal parabolic geometries is valued in the $\mathfrak{g}_{0}$ module $H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$. However, Borel-Weil-Bott theorem only permits computation of the complex analog $H^{2}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}}),$ a $\mathfrak{g}_{-}^{\mathbb{C}}$ module. The following theorem allows us to compare the two modules.
@@
Suppose $\mathfrak{g}$ is noncomplex simple. There is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega\in H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ contained in $V_{\beta,\gamma,\zeta}$ exactly when there is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector $\widetilde{\Omega}\in H^{2}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}})$ of the form
@@
is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector. By Lemma 3.5, there is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ in $V_{\beta,\gamma,\zeta}.$
Now suppose that there is a lowest weight vector $\Omega\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ in $V_{\beta,\gamma,\zeta}$. Let $V\leq H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ be a $\mathfrak{g}_{0}$ irreducible subrepresentation containing $\Omega$. By Proposition 2.8,
@@
Let $\widetilde{\Omega}\in V^{\mathbb{C}}$ be a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector. By the Borel-Weil-Bott theorem, $\widetilde{\Omega}=(\eta_{\widetilde{\beta}})_{\flat}\wedge(\eta_{\widetilde{\gamma}})_{\flat}\otimes\eta_{\widetilde{\zeta}}$ for some roots $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ of $\mathfrak{g}_{0}^{\mathbb{C}}$. By Lemma 3.5, there is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega^{\prime}\in V$ contained in $V_{\beta,\gamma,\zeta}$, where $\beta,\gamma,\zeta$ are the restrictions of $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ to $\mathfrak{a}.$ By Proposition 3.1, $\Omega\in V_{\beta,\gamma,\zeta}.$
Recalling Remark 3.2 and Borel-Weil-Bott theorem, this implies the following.

### v4

@@ 3.4 Curvature Trees and Harmonic Seeds
Given a $\mathfrak{g}$-valued $2$-form $\Omega\in\ker\Delta$, let $K_{\Omega}=\mathrm{Stab}_{G_{0}}(\Omega)$ and let $\mathfrak{k}_{\Omega}\leq\mathfrak{g}_{0}$ be its Lie algebra. A form $\Omega\in(\ker\Delta)_{+}$ is said to have the Kruglikov-The property if
@@
A form $\Omega$ with the Kruglikov-The property is called a harmonic seed if there exists a model geometry $(J_{\Omega},K_{\Omega})$ and an isomorphism of $K_{\Omega}$-representations $\psi:\mathfrak{j}_{\Omega}\to\mathfrak{g}_{-}\oplus\mathfrak{k}_{\Omega}\leq\mathfrak{g}$ such that $\mathfrak{j}_{\Omega}$ is the Lie algebra of $J_{\Omega},$ $\psi|_{\mathfrak{k}_{\Omega}}=1_{\mathfrak{k}_{\Omega}}$ and $J_{\Omega}/K_{\Omega}$ is simply connected. Let $\omega_{J_{\Omega}}$ and $\omega_{P}$ be the Maurer-Cartan forms of the Lie groups $J_{\Omega}$ and $P$. Then the Cartan geometry $\mathcal{G}_{\Omega}:=J_{\Omega}\times_{K_{\Omega}}P$ modeled on $(G,P)$ over $J_{\Omega}/K_{\Omega}$ with Cartan connection given by
@@
is called the curvature tree grown from $\Omega.$ For $j\in J_{\Omega},$ let $L_{j}:\mathcal{G}_{\Omega}\to\mathcal{G}_{\Omega}$ denote left-action by $j.$ This transformation is right $P$-equivariant. Also,
@@
so $J_{\Omega}\leq\mathrm{Aut}(\mathcal{G}_{\Omega},\omega_{\Omega}).$ The following is Theorem 3.4 in [6].
@@
Let $\mathfrak{b}_{-}$ be the nilpotent subalgebra of $\mathfrak{g}$ generated by negative restricted root spaces. If $\Omega\in(\ker\Delta)_{+}$ satisfies the Kruglikov-The property and $\mathrm{im}(\Omega)\subset\mathfrak{b}_{-}$ then $\Omega$ is a harmonic seed.
@@ 4 Lowest Weights
4.1 Lowest Weights in the Harmonic Curvature Module
This section develops results enabling us to compute lowest weight vectors in the module of harmonic curvature forms. With these results in hand, we can prove Theorem 4.11 and Theorem 4.13, which are useful for constructing harmonic seeds and compact quotients of curvature trees, respectively. Let $\mathfrak{g}$ be real semisimple. We keep in mind the identification $H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})\cong\ker\Delta\leq C^{2}(\mathfrak{g}_{-},\mathfrak{g})$. Given $\beta,\gamma\in\hat{\Delta}^{+}(\mathfrak{p}_{+})$ and $\zeta\in\hat{\Delta}$, define a subspace $\mathfrak{g}_{\beta,\gamma,\zeta}:=(\mathfrak{g}_{\beta})_{\flat}\wedge(\mathfrak{g}_{\gamma})_{\flat}\otimes\mathfrak{g}_{\zeta}\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$.
@@ 4.1 Lowest Weights in the Harmonic Curvature Module
Let $V\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$ be a $\mathfrak{g}_{0}$-irreducible representation. Suppose there is a $\mathfrak{g}_{0}$ lowest weight vector $v\in V$ contained in $\mathfrak{g}_{\beta,\gamma,\zeta}.$ Then every $\mathfrak{g}_{0}$ lowest weight vector in $V$ is contained in $\mathfrak{g}_{\beta,\gamma,\zeta}$.
@@
We claim that every lowest weight vector in $V$ is in the $\mathfrak{z}(\mathfrak{g}_{0})$ module generated by $v,$ which implies the result because $\mathfrak{z}(\mathfrak{g}_{0})$ is contained in the Cartan subalgebra of $\mathfrak{g}.$ Let $w\in V$ be another $\mathfrak{g}_{0}$ lowest weight vector. Let $\mathfrak{g}_{0}^{\geq 0}:=\mathfrak{z}(\mathfrak{g}_{0})\oplus\bigoplus_{\alpha\in\hat{\Delta}^{+}(\mathfrak{g}_{0})}\mathfrak{g}_{\alpha}.$ Since $v$ is a lowest weight element, an inductive argument shows that $V$, the $\mathfrak{g}_{0}$ module generated by $v$, is in fact the $\mathfrak{g}_{0}^{\geq 0}$ module generated by $v$. The lowest restricted weight of an irreducible representation is unique, so $v$ and $w$ both have the same restricted weight. Because $w$ is in the $\mathfrak{g}_{0}^{\geq 0}$ module generated by $v$ but has the same restricted weight, it must be in the $\mathfrak{z}(\mathfrak{g}_{0})$ module generated by $v.$
@@
Vectors in $\mathfrak{g}_{\beta,\gamma,\zeta}$ have weight $\beta+\gamma+\zeta.$
Let $\iota_{\mathfrak{a}}:\mathfrak{h}^{*}\to\mathfrak{a}^{*}$ be the restriction map.
Proposition 4.3.
Let $\mathfrak{g}\curvearrowright V$ be a representation with complexification $\mathfrak{g}^{\mathbb{C}}\curvearrowright V^{\mathbb{C}}$. Let $\lambda\in\mathfrak{a}^{*}.$ Then $\bigoplus_{\widetilde{\lambda}\in\iota_{\mathfrak{a}}^{-1}(\lambda)}(V^{\mathbb{C}})_{\widetilde{\lambda}}=(V_{\lambda})^{\mathbb{C}}.$
Proof.
We immediately have $(V_{\lambda})^{\mathbb{C}}\leq(V^{\mathbb{C}})_{\lambda}.$ Since $\bigoplus_{\lambda\in\mathfrak{a}^{*}}(V_{\lambda})^{\mathbb{C}}=V^{\mathbb{C}}=\bigoplus_{\lambda\in\mathfrak{a}^{*}}(V^{\mathbb{C}})_{\lambda}$, none of these inclusions can be proper, and so $V_{\lambda}^{\mathbb{C}}=(V^{\mathbb{C}})_{\lambda}.$
On the other hand, if $\widetilde{\lambda}|_{\mathfrak{a}}=\lambda$, then $(V^{\mathbb{C}})_{\widetilde{\lambda}}\leq(V^{\mathbb{C}})_{\lambda},$ which implies
$\bigoplus_{\widetilde{\lambda}\in\iota_{\mathfrak{a}}^{-1}(\lambda)}(V^{\mathbb{C}})_{\widetilde{\lambda}}\leq(V^{\mathbb{C}})_{\lambda}.$
Indexing over all $\lambda\in\mathfrak{a}^{*},$ both sides direct sum to to $V^{\mathbb{C}},$ which implies
$\bigoplus_{\widetilde{\lambda}\in\iota_{\mathfrak{a}}^{-1}(\lambda)}(V^{\mathbb{C}})_{\widetilde{\lambda}}=(V^{\mathbb{C}})_{\lambda}=(V_{\lambda})^{\mathbb{C}}.$
Specializing to the adjoint representation of $\mathfrak{g}$ we have $\bigoplus_{\widetilde{\alpha}\in\iota_{\mathfrak{a}}^{-1}(\alpha)}(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\alpha}}=(\mathfrak{g}_{\alpha})^{\mathbb{C}}.$ In particular, if $\widetilde{\alpha}|_{\mathfrak{a}}=\alpha$ then $(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\alpha}}\leq(\mathfrak{g}^{\alpha})^{\mathbb{C}}.$
@@
Using the identifications $\mathfrak{p}^{+}\cong\mathfrak{g}_{-}^{*}$ and $\mathfrak{p}_{+}^{\mathbb{C}}\cong(\mathfrak{g}_{-}^{\mathbb{C}})^{*}$ taking $\eta\mapsto\eta_{\flat_{\mathbb{R}}}$ and $\eta\mapsto\eta_{\flat_{\mathbb{C}}}$, the inclusion $\mathfrak{p}_{+}\to\mathfrak{p}_{+}^{\mathbb{C}}$ becomes $\mathbb{C}$-linear extension $\mathfrak{g}_{-}^{*}\to(\mathfrak{g}^{\mathbb{C}}_{-})^{*}.$ Thus the inclusion map $C_{2}^{\mathbb{R}}(\mathfrak{p}_{+},\mathfrak{g})\to C_{2}^{\mathbb{C}}(\mathfrak{p}_{+}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}})$ taking $\eta_{1}\wedge_{\mathbb{R}}\eta_{2}\otimes_{\mathbb{R}}\eta_{3}\mapsto\eta_{1}\wedge_{\mathbb{C}}\eta_{2}\otimes_{\mathbb{C}}\eta_{3}$ is identified with the extension map $C^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})\to C^{2}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}})$ taking $(\eta_{1})_{\flat_{\mathbb{R}}}\wedge_{\mathbb{R}}(\eta_{2})_{\flat_{\mathbb{R}}}\otimes_{\mathbb{R}}\eta_{3}\mapsto(\eta_{1})_{\flat_{\mathbb{C}}}\wedge_{\mathbb{C}}(\eta_{2})_{\flat_{\mathbb{C}}}\otimes_{\mathbb{C}}\eta_{3}$.
Corollary 4.5.
Let $\widetilde{\beta},\widetilde{\gamma}\in\Delta(\mathfrak{p}_{+}^{\mathbb{C}})$ and $\widetilde{\zeta}\in\Delta(\mathfrak{g}^{\mathbb{C}}),$ and let $\beta,\gamma\in\hat{\Delta}(\mathfrak{p}^{+})$ and $\zeta\in\hat{\Delta}(\mathfrak{g})$ be their restrictions to $\mathfrak{a}.$ Then $(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}\leq(\mathfrak{g}_{\beta,\gamma,\zeta})^{\mathbb{C}}$.
@@
From the proposition we know that
$(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\beta}}\wedge(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\gamma}}\otimes(\mathfrak{g}^{\mathbb{C}})_{\widetilde{\zeta}}\leq(\mathfrak{g}_{\alpha})^{\mathbb{C}}\wedge(\mathfrak{g}_{\beta})^{\mathbb{C}}\otimes(\mathfrak{g}_{\zeta})^{\mathbb{C}}=(\mathfrak{g}_{\alpha}\wedge\mathfrak{g}_{\beta}\otimes\mathfrak{g}_{\zeta})^{\mathbb{C}}.$
The result follows from applying the identifications of Remark 4.4. ∎
@@
Suppose $V\leq\bigwedge^{2}(\mathfrak{g}_{-})^{*}\otimes\mathfrak{g}$ is a $\mathfrak{g}_{0}$ subrepresentation, and $\widetilde{\Omega}\in V^{\mathbb{C}}$ is a $\mathfrak{g}_{0}^{\mathbb{C}}$ weight vector such that $\widetilde{\Omega}\in\mathfrak{g}_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}$. Then the real and imaginary parts of $\widetilde{\Omega}$ in $V$, if nonzero, are $\mathfrak{g}_{0}$ weight vectors in $\mathfrak{g}_{\beta,\gamma,\zeta}$, where $\beta,\gamma,\zeta$ are the restrictions of $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ to $\mathfrak{a}.$ In addition, if $\widetilde{\Omega}$ is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector, then its real and imaginary parts, if nonzero, are $\mathfrak{g}_{0}$ lowest weight vectors.
@@
Without loss of generality, we prove the result for the real part. Define $\Omega,\Omega^{\prime}\in V$ as the unique vectors such that $\tilde{\Omega}=\Omega+i\Omega^{\prime}.$ Since $\tilde{\Omega}\in\mathfrak{g}_{\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}}\leq(\mathfrak{g}_{\beta,\gamma,\zeta})^{\mathbb{C}},$ we have $\Omega,\Omega^{\prime}\in\mathfrak{g}_{\beta,\gamma,\zeta}.$ By assumption, $\Omega\neq 0$. Because $\tilde{\Omega}$ is a weight vector for $\mathfrak{g}_{0}^{\mathbb{C}}$ and $\mathfrak{a}\leq\mathfrak{h}_{0}$, $\widetilde{\Omega}$ scales by real values under the adjoint action of $\mathfrak{a}.$ Therefore $\Omega$ is a weight vector for $\mathfrak{g}_{0}.$
Now suppose $\widetilde{\Omega}$ is a lowest weight. Let $\alpha$ be some negative $\mathfrak{g}_{0}$ restricted root. By Proposition 4.3,
$((\mathfrak{g}_{0})_{\alpha})^{\mathbb{C}}=\bigoplus_{\widetilde{\alpha}\in\iota_{\mathfrak{a}}^{-1}(\alpha)}(\mathfrak{g}_{0}^{\mathbb{C}})_{\widetilde{\alpha}}$,
a direct sum of negative rootspaces of $\mathfrak{g}_{0}^{\mathbb{C}}.$ These rootspaces annihilate $\tilde{\Omega}$ by assumption, so $(\mathfrak{g}_{0})_{\alpha}$ annihilates $\tilde{\Omega}$, so $(\mathfrak{g}_{0})_{\alpha}$ annihilates $\Omega.$ It follows that $\Omega$ is a lowest weight vector. ∎
As discussed in Section 3.1, harmonic curvature of regular, normal parabolic geometries is valued in the $\mathfrak{g}_{0}$ module $H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$. However, Kostant’s Borel-Weil-Bott facilitates computation of the complex analog $H^{2}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}}),$ a $\mathfrak{g}_{0}^{\mathbb{C}}$ module. The following theorem allows us to compare the two modules.
@@
Suppose $\mathfrak{g}$ is noncomplex simple. There is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega\in H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ contained in $\mathfrak{g}_{\beta,\gamma,\zeta}$ exactly when there is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector $\widetilde{\Omega}\in H^{2}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},\mathfrak{g}^{\mathbb{C}})$ of the form
@@
is a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector. By Lemma 4.6, there is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ in $\mathfrak{g}_{\beta,\gamma,\zeta}.$
Now suppose that there is a lowest weight vector $\Omega\in H^{2}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ in $\mathfrak{g}_{\beta,\gamma,\zeta}$. Let $V\leq H_{\mathbb{R}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ be a $\mathfrak{g}_{0}$ irreducible subrepresentation containing $\Omega$. By Proposition 3.4(b),
@@
Let $\widetilde{\Omega}\in V^{\mathbb{C}}$ be a $\mathfrak{g}_{0}^{\mathbb{C}}$ lowest weight vector. By the Kostant’s Borel-Weil-Bott, $\widetilde{\Omega}=(\eta_{\widetilde{\beta}})_{\flat}\wedge(\eta_{\widetilde{\gamma}})_{\flat}\otimes\eta_{\widetilde{\zeta}}$ for some roots $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ of $\mathfrak{g}_{0}^{\mathbb{C}}$. By Lemma 4.6, there is a $\mathfrak{g}_{0}$ lowest weight vector $\Omega^{\prime}\in V$ contained in $\mathfrak{g}_{\beta,\gamma,\zeta}$, where $\beta,\gamma,\zeta$ are the restrictions of $\widetilde{\beta},\widetilde{\gamma},\widetilde{\zeta}$ to $\mathfrak{a}.$ By Proposition 4.1, $\Omega\in\mathfrak{g}_{\beta,\gamma,\zeta}.$
Recalling Remark 4.2 and Kostant’s Borel-Weil-Bott, this implies the following.